8. Show that the real valued function defined by f(x) = x 3 − 3x 2 + 3x + 1 increases for all x ∈ R


1
Expert's answer
2021-04-28T16:36:30-0400

Let us show that the real valued function defined by "f(x) = x^3 \u2212 3x^2 + 3x + 1" increases for all "x ∈ \mathbb R." For this let us find the derivative: "f'(x) = 3x^2 \u2212 6x + 3." Taking into account that "f'(x) = 3x^2 \u2212 6x + 3= 3(x^2 \u2212 2x + 1)=3(x-1)^2>0" for all "x ∈ \mathbb R\setminus\{1\}" and "f'(1)=0", we conclude that the function "f" increases for all "x ∈ \mathbb R."